# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v4

@@ 6 Discussion
The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping. Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one-to-one relation between the two. In the case of a network structure, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy, $U$, is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale, $\tau$, tantamount to the inverse temperature, $\beta$, in statistical physics. Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau$ [3]. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given the value of the external control parameter, $r$.
Notably, we have computed the thermodynamic efficiency at a specific resolution scale, $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e. from star graph to amorphous. The operational $\tau$-scale employed roughly identifies the scale where modules emerge in the network. For this reason, we could gauge the phase transition employing the network modularity, $Q$, as an order parameter. While the transition to scale-invariant permits the identification of the network resolution scale, $\tau$, where the thermodynamic efficiency needs to be measured to observe a peak, the corresponding transition value of the control parameter, $r$, does not always coincide with the global maximum in the efficiency. We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of only the Portegys algorithm. Given a pre-trained tree network based upon an unknown dataset, it might be possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the peaks of the pattern-matching and thermodynamic efficiency on the phase plane without performing any test phase. As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, the peak of the pattern-matching efficiency might be at the phase transition boundary providing higher accuracy at the cost of a slower pattern search. Interestingly, our analysis of hierarchical tree networks could be applied, if further developed, to other topics in artificial intelligence. We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing. In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals. Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years. In the neuromorphic paradigm, computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48]. Tree-like structures are currently physically implemented in materio. We mention tree structures produced at the critical point of a percolation transition [49, 50] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 51], but also fractal trees produced in the dielectric breakdown [52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
@@
This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program ”Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No. 861153, as well as by the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).
@@ Acknowledgments
Supplemental Data
The Supplementary Material includes figures referenced in the main text but omitted for clarity.
@@
The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.

### v5

@@ IV Discussion
The matter is what makes a certain structure better suited in terms of an objective, such as accuracy or efficiency, to a given task. Equivalently and more generally said, in terms of a complex input-output mapping. When dealing with artificial neural networks, we can either enlarge the number of parameters or adjust their values to achieve more complicated functions, even though neither implies a one-to-one relationship with better generalization performance. In the case of our network structures, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure a weighted average of the parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy, $U$, is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale, $\tau$, tantamount to the inverse temperature, $\beta$, in statistical physics. Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau=\frac{1}{2}\frac{d_{s}}{\tau}$ [35] in a limited $\tau$ range due to the finite size of network and the frequency cut-off that we discuss in the following. Thus, $U$ is related to the specific heat plateau, $C_{0}$, and, importantly, to the spectral dimension [41], $d_{s}$, which governs the spectral density. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, or directly adjust the parameter values. Still, we can govern the distribution of frequencies with the control parameter, $r$, and let the structure self-organize.
We have computed the thermodynamic efficiency at a specific resolution scale, $\tau=10$, identified by inspecting the specific heat, $C=-dS/d\log\tau$. At this $\tau$-scale, we could monitor the transition from the trivial structure, a single specific heat peak, to the emergence of complex amorphous structures, where the specific heat shows a plateau. The plateau reflects the growth of an approximately scale-invariant regime where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an “ultra-violet” cut-off, delimiting from above the relevant frequencies of the networks across the transition to scale-invariance. At the same time, it scans the scale where modules emerge in the (loop-less) network. For this reason, we could gauge the phase transition to the amorphous regime with the network modularity, $Q$, as an order parameter. Finally, this resolution $\tau$-scale is independent of the control parameter, the dataset, and the doubling of the number of nodes (see Supplemental Material [63]).
Highly efficient networks were always found in the amorphous regime. From a functional point of view, we mention that several works in the literature [67, 11, 68] suggest that a system benefits from the vicinity to a phase transition critical point when it confronts environments of high complexity, while it does not benefit from it when the environmental complexity is lower. This partly aligns with our findings on the four datasets of increasing complexity. One key aspect of being close to a critical point is to take advantage of the functional benefits that a power law density related to some quantity of interest has to offer. In our case, for all datasets, we saw higher performance in terms of pattern-matching efficiency in the amorphous region of the phase plane, characterized by a power law spectral density.
Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star graph and the scale-invariant regime. On the contrary, lower complexity datasets find their highly efficient structures on the opposite side of the amorphous phase plane region.
Then what is the reason for a network embedding high dimensional patterns to be in one or the other location of the phase plane? By measuring the optimally efficient networks’ spectral dimension [39, 41], $d_{s}$, almost continuously changing in the amorphous regime, we could relate the dimensionality of the network to the complexity of the dataset it structured. We observed higher spectral dimension, reflecting large-scale connectedness, for networks mapping high-complexity datasets. On the contrary, easier datasets were mapped into more complex branching structures with a lower spectral dimension. Thus, importantly, we provide an alternative way to measure the complexity of a dataset: by mapping it into a tree network and then measuring its spectral dimension.
Returning to the starting question of this manuscript—why a certain structure emerges in place of another—the answer suggested by our findings is: to manage the complexity of the environment it is adapting to.
The generalization of this possible answer to systems other than the Portegys algorithm remains purely speculative and beyond the scope of this work. Still, the methods presented can in principle provide diagnostics to other tree-based approaches to classification such as hierarchical clustering.
Suppose we are given a pre-trained tree network mapping an unknown dataset, we can estimate the complexity of the dataset that was used for training, along with the ability of the tree structure to cope with that complexity without performing any test phase. Suppose the trained network “struggles” in the representation of the dataset. In such a case, the spectral dimension is expected to be high, and the network is expected to exhibit higher connectedness. On the contrary, a network mapping a low-complexity dataset should exhibit branched structures and low spectral dimension.
In the reverse direction, knowing the complexity of a dataset, we could initialize the search for the optimal control parameter, $r$, closer to values typically presenting one of the two transitions exhibited by the algorithm, reducing the number of simulations required to find the most efficient network structure.
Interestingly, developments of our analysis and methods on hierarchical tree networks could be applied to other topics in artificial intelligence, neuroscience, and neuromorphic computing.
We find it particularly suited to the context of biologically inspired [16, 17, 15] dendritic learning [20, 21, 22, 23], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for parameter-efficient learning in artificial neural network [20, 21, 22, 23].
Neurons can be modeled as the amorphous tree structures discussed in this work. It would be interesting to experimentally relate the dendritic structure, specifically its spectral dimension, to the capability of the neuron to deal with environmental stimuli and complicated input-output mapping of signals [15, 16, 17, 18, 19], or to produce signals with richer diversity [74, 71, 75]. For instance, a neuron capable of dealing with the complexity of the stimulus it is subject to would be expected to exhibit traits with low spectral dimensionality and complex branching.
Moreover, potential applications include the topic of neuromorphic computing [44, 76, 77] currently looking for a possible solution to the ever-increasing computational power demand of recent years. Neuromorphic computing processors are designed to co-localize computation and memory on the physical substrate and emulate the brain learning capabilities in materia. Complete neural networks with dendritic branching have been fabricated with memristive devices and have shown significant power efficiency [24]. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [78]. Analogously to biological neural systems, structurally efficient routing may play a key role for in-materia computing processors fabricated with self-assembled materials [79, 80]. Emergent tree-like structures are currently physically fabricated in emerging materials representing potential candidates for neuromorphic computing. We mention tree structures produced at the critical point of a percolation transition for randomly dispersed nanoparticles [25, 81], or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [25, 82], and fractal trees produced in the dielectric breakdown [83]. Thus, their information processing capabilities could be investigated and enhanced based on this work and future developments.
V Conclusions
We summarize the contributions included in this work.
(i) We proposed the Portegys algorithm [27, 28] as an alternative process to produce hierarchies for statistical physics analysis, integrating the functional aspect of a machine-learning algorithm.
(ii) Multiple phase transitions are identified while the algorithm searches for an efficient internal representation of the dataset.
(iii) Scale-invariance, i.e. power-law Laplacian spectral density [36], is a key feature to constructing efficient tree-networks, capable of combining fast information flow and sufficiently rich internal representation of information. Our analysis involved task-agnostic thermodynamic-like efficiency [1], based on general statistical physics arguments, and task-related pattern-matching efficiency, quantifying the trade-off between accuracy and resources used to achieve the task.
(iv) Studying efficient tree structures mapping the structured disorder of a dataset, we showed that the emergence of complex trees is directly influenced by the complexity of the environmental conditions. Namely, the spectral dimension [39, 40, 41], proportional to the exponent of the power-law spectral density, directly encodes the complexity of the disordered environment.
(v) Therefore, we proposed two novel operational measurements to estimate the dataset complexity: by measuring the spectral dimension of the highly efficient network mapping the dataset; or, equivalently, by measuring the specific heat plateau [34, 35], which reflects the spectral dilation symmetry and the network resolution where information is integrated at a constant rate.
@@ V Conclusions
This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program “Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No. 861153, as well as by the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).
@@ Acknowledgments
The original contributions presented in the study are included in the article/supplemental material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.